OpenStudy (anonymous):

Find the surface area of the ball. Leave the answer in terms of PIE d = 24 cm

OpenStudy (anonymous):

Radius = 24/2 = 12cm Surface area = 4 x pi x 12^2 = 576pi cm^2

OpenStudy (anonymous):

THANKSSS! HAHA XD :)

OpenStudy (anonymous):

Xoxoxoxo <3

OpenStudy (anonymous):

You're welcome!! Xo! ;)